OpenStudy (anonymous):

What is the 11th term in the sequence: an = 6n + 13? 79 - 79 -53 53

jimthompson5910 (jim_thompson5910):

Replace every n with 11 and evaluate

jimthompson5910 (jim_thompson5910):

an = 6n + 13 a11 = 6(11) + 13 a11 = ??

OpenStudy (jackthegreatest):

lets put it short, @brittney_fraley the anser is 79

OpenStudy (anonymous):

thanks yall

OpenStudy (jackthegreatest):

medal?